To solve the initial value problem using the method of Laplace transforms:
y''+16y=8cos4t; y(0)=0; y'(0)=8
1
Expert's answer
2020-01-31T12:33:54-0500

"y''+16y'=8\cos4t\\
y(0)=0, y'(0)=8\\"

Laplace transforms is

"y(t)\mapsto Y\\
y''(t)\mapsto p^2Y-py(0)-y'(0)=\\
=p^2Y-p•0-8=p^2Y-8 \\
\cos4t\mapsto\frac{p}{p^2+16}"

then

"p^2Y-8+16Y=\frac{8p}{p^2+16}\\
(p^2+16)Y=\frac{8p}{p^2+16}+8\\
Y=\frac{8p}{(p^2+16)^2}+\frac{8}{p^2+16}"

Then

"\frac{8p}{(p^2+16)^2}=\\
=\frac{2p•4}{(p^2+4^2)^2}\mapsto t\sin 4t\\"

"\frac{8}{p^2+16}=\frac{2•4}{p^2+4^2}\mapsto 2\sin 4t"

Solution is

"y=t\sin4t+2\sin4t\\
y=(t+2)\sin4t"